The perimeter of a rectangle is 260 m. If the length is 5 m less than twice the width, find the dimensions of the rectangle.
What is the length?
What is the width?

Answer by macston(5194)

W=width; L=length=2W-5; Perimeter=P=2(L+W)=260 m
P=2(L+W) Substitute for L
260 m=2((2W-5)+W)
260 m=2(3w-5) Divide each side by 2
130 m=3W-5 Add 5 to each side
135 m=3W Divide each side by 3
45 m=W ANSWER 1: Width is 45 meters.
L=2W-5=2(45)-5=90-5=85 m ANSWER 2: Length is 85 meters
CHECK:
P=260 m=2(L+W)=2(85 m+45 m)
260 m=2(130 m)
260 m=260 m